# Parallel Scan on Ascend AI Accelerators

### v1

@@ Abstract.
Our single-core results demonstrate substantial performance improvements, with speedups ranging from 5×5$\times$5 × to 9.6×9.6$\times$9.6 × compared to vector-only implementations for sufficiently large input lengths. Additionally, we present a multi-core scan algorithm that fully utilizes both the cube and vector units of Ascend, reaching up to 37.5% of the theoretical memory bandwidth. Furthermore, our radix sort implementation, which utilizes matrix multiplications for its parallel splits, showcases the potential of matrix engines to enhance complex operations, offering up to $\times3$ speedup over the baseline.
prefix sum, scan, matrix multiplication, matrix engines, tensor cores, Ascend, AI accelerators
copyright: Nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Unpublished manuscript; May 2025; N/Aisbn: 978-1-4503-XXXX-X/18/06ccs: Theory of computation Parallel algorithms
1. Introduction
@@
Parallel scan is a fundamental parallel computing paradigm with many applications (Blelloch, 1990; Lakshmivarahan and Dhall, 1994). Due to its importance, parallel scan has been studied in several models of computation, including the circuit (work and depth) and the Parallel Random-Access Machine (PRAM) model (Blelloch, 1996). In the circuit model, the depth and size trade-offs for parallel optimal prefix circuits are well-understood for binary operations (Snir, 1986; Zhu et al., 2006), but other models are yet to be explored, especially in the heterogeneous computing domain (Dakkak et al., 2019; Zouzias and McColl, 2023).
@@
Despite many attempts, there is a large gap between abstract parallel machine models and the current state-of-the-art accelerators that contain heterogeneous computing units, see(Matias, 1997) and (Arge et al., 2008; Sitchinava and Weichert, 2013; Koike and Sadakane, 2014). For example, hardware vendors have manufactured accelerators with specialized compute engines known as matrix engines or tensor core units. A list of such specialized hardware units includes Google’s TPUs (Jouppi and et al., 2017, 2021), Nvidia’s Tensor Cores (NVIDIA Authors, 2017), AMD’s Matrix Cores (AMD, 2022) and Huawei’s Ascend Cube Unit (Liao et al., 2021, 2019) to name a few. Moreover, in the CPU domain, examples of such units (or extensions) are ARM’s Scalable Matrix Extension (SME) (ARM, 2024), Intel’s Advanced Matrix Extensions (AMX) (Intel Corporation, 2022) and IBM’s POWER10 Matrix-Multiply Assist (MMA) (Starke et al., 2021). Therefore, today’s high-performance processors contain matrix engines that allow efficient multiplication of constant-sized matrices, making ideas from the early 1980s (Kung and Leiserson, 1978; Snyder et al., 1985) a reality and initiating fruitful debates within the community (Domke et al., 2021).
@@
Given the presence of matrix engines in computing products, a model of computation was recently proposed to capture matrix multiplication accelerators called the Tensor Core Unit (TCU) model (Chowdhury et al., 2020, 2021). The authors of (Chowdhury et al., 2021) initiated the study of TCU algorithms and revisited classical paradigms in the TCU model. From the perspective of hardware-aware algorithmic implementations, the seminal paper of (Dakkak et al., 2019) proposed and implemented parallel scan (and reduction) kernels tailored to GPU accelerators. We drew inspiration from both these works (Dakkak et al., 2019; Chowdhury et al., 2021) to advance the study of parallel scans on matrix multiplication accelerators. As a case study and an evaluation environment, we use Huawei’s Ascend AI accelerators and the Compute Architecture for Neural Networks (CANN) software ecosystem of Ascend to evaluate our proposals (Liang, 2020).
@@
  • We evaluated the proposed scan algorithms that use a single cube and vector units, demonstrating a 5×5$\times$5 × up to 9.6×9.6$\times$9.6 × speed-up compared to the baseline scan (vector unit only) for sufficiently large input lengths.
  • We present a multi-core scan (MCScan) algorithm that fully utilizes both the cube and vector units of the Ascend 910B4 accelerator, reaching up to 37.5% of the theoretical memory bandwidth. MCScan achieves a 15.2×15.2$\times$15.2 × speedup compared to the single cube algorithm when it uses all available (20) cube cores and vector cores.
@@
  • Lastly, we implement radix sort, whose parallel splits take advantage of the cube units. The radix sort implementation provides up to $\times3$ speed-up over baseline sorting that does not utilize the cube units.
@@
Second, our radix sort implementation yields an intriguing result: in practice, multiple small dense matrix multiplications can be leveraged to improve the end-to-end performance of parallel sorting. Although the algorithmic ideas and techniques underlying this result are well-established (Blelloch et al., 1991), we believe that it paves the way for interesting research directions in the future. As an example, we pose the following question. Is it possible to utilize the multiple-add capabilities of the matrix multiplications units to improve further the performance of parallel sorting? A similar question of ”Can we sort with matrix multiplications?” was raised during the presentation of (Chowdhury et al., 2020). Note that although the parallel splits of our radix sort implementation heavily use matrix multiplications to perform scans, they do not use the multiply-add capabilities of the matrix multiplication unit.
2. Background
In this section, we give an overview of related work in computing parallel scans using accelerators, focusing mostly on matrix multiplication accelerators. We also present the Tensor Core Unit (TCU) model of computation, and we give a summary of a selected list of scan-based primitives. The (inclusive) prefix sum or scan of a sequence is a new sequence where each term at index i is the cumulative sum of all terms from the beginning of the original sequence up to index i.
2.1. Scan Strategies on Accelerators
Numerous parallel scan implementations and libraries have been proposed in the literature (Dotsenko et al., 2008; Yan et al., 2013; Hwu et al., 2023a; Sengupta et al., 2007; Baxter, 2016; Bell and Hoberock, 2012; NVIDIA, 2023; Merrill and Garland, 2016). Here, we discuss the most relevant work focusing on scan implementations targeting accelerators, most notably Graphics Processing Units (GPUs). Horn was one of the first to implement parallel scans in GPUs (Pharr and Fernando, 2005, Chapter 36), followed up with several improvements.
In a nutshell, GPU scan implementations primarily follow a two-level hierarchical approach where the highest level of the hierarchy is the block level. An efficient scan implementation follows one of these scan strategies: Scan-Scan-Add (SSA), Reduce-Scan-Scan (RSS), or Stream-scan according to the state-of-the-art decouple lookback scan approach of (Merrill and Garland, 2016). Scan-Scan-Add (SSA) means that the (local) scans are initially computed per block. Second, the values of the largest index of each block are collected and scanned. Third, the collected per-block scan values are broadcast-added to their corresponding blocks. Similarly, the RSS approach makes a block-level reduction first, followed by a scan of the block-level reductions and a final scan of each block. We refer the interested reader to (Merrill and Garland, 2016, Section 3) for a more detailed discussion.
@@ 2.1. Scan Strategies on Accelerators
StreamScan and decouple look-back strategies access only ${}2N$ memory elements but need to efficiently handle the sequential data dependency of the scan computation using adjacent (block) synchronization, see  (Hwu et al., 2023b, Chapter 11.7). StreamScan is a single-pass approach in which each thread block is assigned a tile of input, and a serial dependency between the blocks exists (Yan et al., 2013). The critical feature of StreamScan is that it requires synchronization between adjacent blocks only (without global block-level synchronization). The decoupled look-back strategy of (Merrill and Garland, 2016) aims to alleviate the drawbacks of the serial dependency of StreamScan by performing redundant work to “dissociate” local computation from the latency of global prefix propagation. Both StreamScan and decouple look-back strategies require only ${}2N$ data movement to global memory: $N$ input elements are read, $N$ output elements are written.
2.2. Scan using Matrix Multiplications
The study of accelerating prefix sum (and reduction) operations using matrix multiplication units was first initiated in the seminal paper of (Dakkak et al., 2019). The authors of (Dakkak et al., 2019) designed scan algorithms for the GPU architecture by providing highly optimized CUDA kernels and, hence, use the terms of warp/block/grid of the CUDA programming model. They proposed a warp-level scan algorithm (Dakkak et al., 2019, Algorithm 6), and a block-level scan algorithm (Dakkak et al., 2019, Algorithm 7). Moreover, they mentioned that the device/grid level algorithm is based on a textbook approach (Hwu et al., 2023a).
A follow-up work presented a parallel scan algorithm in the TCU model of computation where only matrix multiplication operations are required and, more importantly, the depth of the computation is logarithmic in the input size (Zouzias and McColl, 2023). Although the main algorithm of (Zouzias and McColl, 2023) has linear work and logarithmic depth/span, its strided memory access patterns are typically hard to translate to efficient memory access operations.
2.3. Tensor Core Unit (TCU) Model
@@
To the best of our knowledge, the Tensor Core Unit (TCU) model is the only model of computation that has been recently proposed to capture matrix multiplication accelerators (Chowdhury et al., 2020). The TCU model is a standard RAM model with an additional circuit, named tensor core unit, that performs matrix multiplication between constant-size matrices. Although the TCU model captures well a single matrix multiplication computational unit of today’s accelerators, it ignores other essential features: the presence of vector processing units and, more critically, the multi-core nature of these accelerators. Since the TCU model considers only a single matrix multiplication unit, it does not allow its users to conduct a work/depth algorithmic analysis (Blelloch, 1996). Due to the above limitations, any algorithmic analysis in the TCU model will not correspond to a realistic execution in Ascend, i.e., ignoring parallelism and the vector units. Nevertheless, we discuss the work/depth asymptotic analysis of the proposed algorithms, assuming the presence of multiple matrix engines and vector units, considering their operations as basic operations.
2.4. Applications of Parallel Scan
@@
Parallel scan has a plethora of applications (Lakshmivarahan and Dhall, 1994; Blelloch, 1990). Here, we restrict our attention to scan applications that enable us to generate efficient computational kernels (operators) that appear in AI workloads. In particular, we have identified that sorting, weighted sampling, masking of tensors, and top-$k$/top-$p$ sampling are essential. All these applications of the parallel scan are well-known, but the observation that top-$p$ sampling can benefit from scan seems to be new (to the best of our knowledge).
3. Ascend AI Accelerators
@@
In this section, we briefly discuss the DaVinci architecture of Ascend accelerators consisting of the cube and vector computing units, mostly following (Liao et al., 2019). Moreover, we discuss the AscendC programming model of Ascend, a recently proposed programming model for Ascend operator development. All the material presented here is available online at https://www.hiascend.com.
3.1. Ascend Hardware
@@
Figure 1 shows the Ascend architecture where the Cube and Vector units are separate cores. In the 910B architecture, data can only be exchanged using global memory and/or L2 cache. This approach dramatically simplifies the implementation logic, but each data transfer between the AIC and AIV cores might be expensive in terms of performance.
@@
Figure 1. Architecture of Ascend 910B Training series accelerators. Each AI core contains one cube and two vector units.

### v2

@@ Abstract
Our single-core results demonstrate substantial performance improvements, with speedups ranging from $5\times$ to $9.6\times$ compared to vector-only implementations for sufficiently large input lengths. Additionally, we present a multi-core scan algorithm that fully utilizes both the cube and vector units of Ascend, reaching up to 74.9% of the memory bandwidth achieved by memory copy. Furthermore, our radix sort implementation, which utilizes matrix multiplications for its parallel splits, showcases the potential of matrix engines to enhance complex operations, offering up to $3.3\times$ speedup over the vector-only baseline.
1 Introduction
@@
Parallel scan is a fundamental parallel computing paradigm with many applications [7, 30]. Due to its importance, parallel scan has been studied in several models of computation, including the circuit (work and depth) and the Parallel Random-Access Machine (PRAM) model [9]. In the circuit model, the depth and size trade-offs for parallel optimal prefix circuits are well-understood for binary operations [46, 52], but other models are yet to be explored, especially in the heterogeneous computing domain [13, 53].
@@
Despite many attempts, there is a large gap between abstract parallel machine models and the current state-of-the-art accelerators that contain heterogeneous computing units, see [37] and [2, 45, 27]. For example, hardware vendors have manufactured accelerators with specialized compute engines known as matrix engines or tensor core units. A list of such specialized hardware units includes Google’s TPUs [24, 25], Nvidia’s Tensor Cores [41], AMD’s Matrix Cores [1] and Huawei’s Ascend Cube Unit [34, 35] to name a few. Moreover, in the CPU domain, examples of such units (or extensions) are ARM’s Scalable Matrix Extension (SME) [3], Intel’s Advanced Matrix Extensions (AMX) [22] and IBM’s POWER10 Matrix-Multiply Assist (MMA) [48]. Therefore, today’s high-performance processors contain matrix engines that allow efficient multiplication of constant-sized matrices, making ideas from the early 1980s [29, 47] a reality and initiating fruitful debates within the community [14].
Given the presence of matrix engines in computing systems, a model of computation was recently proposed to capture matrix multiplication accelerators called the Tensor Core Unit (TCU) model [11, 12]. The authors of [12] initiated the study of TCU algorithms and revisited classical paradigms in the TCU model. The study of accelerating prefix sum (and reduction) operations using matrix multiplication units was first initiated in the seminal paper of [13]. A follow-up work presented a logarithmic-depth parallel scan algorithm in the TCU model of computation, where only matrix multiplication operations are required [53]. Although the main algorithm in [53] offers stronger theoretical guarantees than that of [13], its poor memory access patterns limit its practical performance.
Here, we focus on simpler algorithms inspired by [13] that achieve performance close to hardware peak. Our top-performing multi-core scan kernels combine ideas from [13, Alg. 6], explicit software-managed asynchronous pipelining between matrix and vector cores, and a novel matrix-vector recomputation strategy in the up-sweep scan phase. Furthermore, we demonstrate that our proposed scan kernels can be easily composed to deliver the best-performing kernels for AI workloads on Ascend.
@@
In short, we draw inspiration from both these works [12, 13] to advance the study of parallel scans on matrix multiplication accelerators. As a case study and an evaluation environment, we use Huawei’s Ascend AI accelerators and the Compute Architecture for Neural Networks (CANN) software ecosystem of Ascend to evaluate our proposals [33].
@@
  • We evaluated the proposed scan algorithms that use a single cube and vector units, demonstrating a $5\times$ up to $9.6\times$ speed-up compared to the baseline scan (vector unit only).
  • We present multi-core scan algorithms that fully utilize both the cube and vector units of the Ascend 910B4 accelerator, reaching up to 74.9% of the theoretical memory bandwidth of memory copy. Moreover, the MCScan algorithm (Algorithm 3) attains a $15.2\times$ speedup over the corresponding single-cube variant when utilizing all 20 available AI cores.
@@
  • Lastly, we implement radix sort, whose parallel splits take advantage of the cube units. The radix sort implementation provides up to $3.3\times$ speed-up over baseline sorting that does not utilize the cube units.
@@
Second, our radix sort implementation yields an intriguing result: in practice, multiple small dense matrix multiplications can be leveraged to improve the end-to-end performance of parallel sorting. Although the algorithmic ideas and techniques underlying this result are well-established [10], we believe that it paves the way for interesting research directions in the future. As an example, we pose the following question, see also the presentation of [11]. Is it possible to also utilize the multiple-add capabilities of the matrix multiplications units to improve further the performance of parallel sorting? It is important to note that, despite relying on the matrix multiplications for scan operations during parallel splits, our radix sort does not exploit the multiply-add capability of the matrix unit.
2 Background
In this section, we give an overview of related work in computing parallel scans using accelerators, focusing mostly on matrix multiplication accelerators. We also present the Tensor Core Unit (TCU) model of computation, and we give a summary of a selected list of scan-based primitives. The inclusive prefix sum of a sequence of elements $x_{1},x_{2},x_{3},\dots$ equipped with a binary associative operator $\odot$ is the sequence $x_{1},x_{1}\odot x_{2},x_{1}\odot x_{2}\odot x_{3},\dots$.
2.1 Scan Strategies on Accelerators
Numerous parallel scan implementations and libraries have been proposed in the literature [15, 50, 20, 44, 4, 5, 40, 38]. Here, we discuss the most relevant work focusing on scan implementations targeting accelerators, most notably Graphics Processing Units (GPUs). Horn was one of the first to implement parallel scans in GPUs [43, Chapter 36], followed up with several improvements.
In a nutshell, GPU scan implementations primarily follow a two-level hierarchical approach where the highest level of the hierarchy is the block level. An efficient scan implementation follows one of these scan strategies: Scan-Scan-Add (SSA), Reduce-Scan-Scan (RSS), or Stream-scan according to the state-of-the-art decouple lookback scan approach of [38]. Scan-Scan-Add (SSA) means that the (local) scans are initially computed per block. Second, the values of the largest index of each block are collected and scanned. Third, the collected per-block scan values are broadcast-added to their corresponding blocks. Similarly, the RSS approach makes a block-level reduction first, followed by a scan of the block-level reductions and a final scan of each block. We refer the interested reader to [38, Section 3] for a more detailed discussion.
@@ 2.1 Scan Strategies on Accelerators
StreamScan and decouple look-back strategies access only 2N~2N memory elements but need to efficiently handle the sequential data dependency of the scan computation using adjacent (block) synchronization, see  [21, Chapter 11.7]. StreamScan is a single-pass approach in which each thread block is assigned a tile of input, and a serial dependency between the blocks exists [50]. The critical feature of StreamScan is that it requires synchronization between adjacent blocks only (without global block-level synchronization). The decoupled look-back strategy of [38] aims to alleviate the drawbacks of the serial dependency of StreamScan by performing redundant work to “dissociate” local computation from the latency of global prefix propagation. Both StreamScan and decouple look-back strategies require only 2N~2N data movement to global memory: $N$ input elements are read, $N$ output elements are written.
To the best of our knowledge, the state-of-the-art scan algorithm on GPUs follows a highly-tuned decoupled look-back and backoff approach, which is briefly explained in [17].
2.2 Tensor Core Unit (TCU) Model
@@
To the best of our knowledge, the Tensor Core Unit (TCU) model is the only model of computation that has been recently proposed to capture matrix multiplication accelerators [11]. The TCU model is a standard RAM model with an additional circuit, named tensor core unit, that performs matrix multiplication between constant-size matrices. Although the TCU model captures well a single matrix multiplication computational unit of today’s accelerators, it ignores other essential features: the presence of vector processing units and, more critically, the multi-core nature of these accelerators. Since the TCU model considers only a single matrix multiplication unit, it does not allow its users to conduct a work/depth algorithmic analysis [9]. Due to the above limitations, any algorithmic analysis in the TCU model will not correspond to a realistic execution in Ascend, i.e., ignoring parallelism and the vector units. Nevertheless, we discuss the work/depth asymptotic analysis of the proposed algorithms, assuming the presence of multiple matrix engines and vector units, considering their operations as basic operations.
2.3 Applications of Parallel Scan
@@
Parallel scan has a plethora of applications [30, 7]. Here, we restrict our attention to scan applications that enable us to generate efficient computational kernels (operators) that appear in AI workloads. In particular, we have identified that sorting, weighted sampling, masking of tensors, and top-kk/top-$p$ sampling are essential. All these applications of the parallel scan are well-known, but the observation that top-$p$ sampling can benefit from scan seems to be new (to the best of our knowledge).
3 Ascend AI Accelerators
@@
In this section, we briefly discuss the DaVinci architecture of Ascend accelerators consisting of the cube and vector computing units, mostly following [35]. Moreover, we discuss the AscendC programming model of Ascend, a recently proposed programming model for Ascend operator development. All the material presented here is available online at https://www.hiascend.com.
3.1 Ascend Hardware
@@
Figure 3.1 shows the Ascend architecture where the Cube and Vector units are separate cores. In the 910B architecture, data can only be exchanged using global memory and/or L2 cache.
@@
Figure 3.1: Architecture of Ascend 910B accelerators. Each AI core contains one cube and two vector units.
4 Scan Algorithms on Ascend
Matrix multiplication is an essential operator in our discussions here; hence, we introduce some linear algebraic notation that will be used throughout the paper. We denote matrices using boldface font and capitals, i.e., $\bm{A},\bm{B},\bm{C}$. We use $s$ to denote the size of square matrices, i.e., $\bm{A}_{s}$, and define $\ell:=s^{2}$. We drop the subscript on the matrices when the dimension is clear from the context. We denote matrix multiplication between $\bm{A}$ and $\bm{B}$ by $\bm{C}:=\bm{A}\ @\ \bm{B}$. We denote by $\bm{U}_{s}$ the upper-triangular all-ones square matrix of size $s$, including ones on the main diagonal. $\bm{L}_{s}$ is the lower triangular all-ones of size $s$. $\bm{L}^{-}_{s}$ is the strictly lower triangular all-ones of size $s$: $\bm{L}_{s}^{-}$ has zeroes on the main diagonal. $\bm{1}_{s}$ denotes the all-ones square matrix of size $s$. We denote one-dimensional arrays using boldface font and lower-case: $\bm{r},\bm{x},\bm{y},\bm{z}$. We frequently partition an array $\bm{x}$ into tiles of length $\ell$, i.e., tiles are contiguous blocks of $\ell$ entries of $\bm{x}$. We note an arbitrary $\ell$-tile of $\bm{x}$ by $\bm{x}_{\ell}$. We also view a tile $\bm{x}_{\ell}$ as a row-major matrix $\bm{A}$ having $s$ rows and $s$ columns (zero-pad if necessary).
In Section 4.2, we present a multi-core scan algorithm (MCScanAlgorithm 3) tailored for the Ascend AI accelerator. A key feature of MCScan is that it utilizes all the available cube and vector cores. MCScan and its extensions are designed for scenarios involving very large one-dimensional arrays.
In Appendix B.1, we build on top of the ScanU and ScanUL1 algorithms and extend them into batched scan variants that operate on multi-dimensional arrays (tensors). We highlight that such extensions are non-trivial, as several issues arise when scheduling many scan operations across multiple cores, including padding and improving load balancing on vector/cube units with a 2:1 ratio, among others.
4.1 Warm-up: single cube scans
Figure 4.1: Data path from an input tile $\bm{x}_{\ell}$ to an output tile $\bm{y}_{\ell}$ of the ScanU (Algorithm 1).
Both algorithms process their tiles using a pipelined execution strategy. The critical path or span of both proposed algorithms is linear on the input length for constant values of $s$, since there is a sequential dependency on the partial sums. Therefore, these kernels are more effective when the input array to be scanned has a relatively short length. Additionally, designing and implementing scan algorithms that use a single cube core is a building block for extending these ideas to a multi-core scenario, as seen in Sections B.1 and 4.2.
Figure 4.1 depicts the data movements and the memory view of ScanU. The input vector $\bm{x}$ is shown in blue. A tile $\bm{x}_{\ell}$ of $\bm{x}$ and $\bm{U}_{s}$ are loaded into the cube unit where the matrix multiplication occurs. The matrix multiplication result is written to global memory. The vector unit reads the cube unit output tile and propagates the prefix sum in place. The whole process, which consists of memory transfers and cube/vector operations, is pipelined over the input tiles using AscendC software pipeline capability.
1:procedure ScanU($\bm{x}$, $s$)
2:  Let $\bm{y}$ be the output array
3:  $\textit{partial}\leftarrow 0$ $\triangleright$ Accumulation value (Vector Unit)
4:  Load $\bm{U}_{s}$ in $\mathtt{L0B}$
5:  for each $s^{2}$-tile of $\bm{x}$: $\bm{x}_{\ell}$ do $\triangleright$ Pipelined exec.
6:   Load $\bm{x}_{\ell}$ to $\mathtt{L0A}$ $\triangleright$ Cube unit
7:   $\bm{C}\leftarrow\bm{A}_{s}\ @\ \bm{U}_{s}$ $\triangleright$ acc. off, free $\mathtt{L0A}$
8:   Copy $\bm{C}$ from $\mathtt{L0C}$ to $\bm{y}_{\ell}$ in $\mathtt{GM}$
9:   Vector unit waits tile from Cube unit
11:   for each $s$-tile of $\bm{y}_{\ell}$: $\bm{y}_{s}$ do
12:     $\bm{y}_{s}\leftarrow\bm{y}_{s}+\textit{partial}$
13:     $\textit{partial}\leftarrow$ last entry of $\bm{y}_{s}$
14:   Copy $\bm{y}_{\ell}$ from $\mathtt{UB}$ to $\mathtt{GM}$
15:  Return $\bm{y}$
Algorithm 1 Scan Cube-Vector (ScanU)
The second algorithm, ScanUL1 (Algorithm 2), is an Ascend adaptation of [13, Algorithm 6] and is based on a matrix identity that expresses the scan of an array $\bm{z}$ of length $\ell$ using matrix operations. View $\bm{z}$ as a square row-major matrix $\bm{A}$ of size $s=\lceil\sqrt{\ell}\rceil$ (pad with zeroes if needed). Given $\bm{z}$, the inclusive scan of $\bm{z}$ ($\texttt{scan}(\bm{z})$) can be computed as:
(4.1) $\texttt{scan}(\bm{z})=\bm{A}_{s}\ @\ \bm{U}_{s}+\bm{L}^{-}_{s}\ @\ \bm{A}_{s}\ @\ \bm{1}_{s},$
ignoring any padded values. Equation 4.1 first appeared in [13]. ScanUL1 uses Equation 4.1 to scan each consecutive tile of size $\ell$ (Lines 6-12 of Algorithm 2), and then propagates the last value of the partial sums sequentially. In a high-level, for each tile of size $\ell$, the cube unit evaluates Equation 4.1 with the following sequence of matrix operations:
$\bm{C}_{1}$ $=\bm{A}_{s}\ @\ \bm{1}_{s}$
$\bm{C}_{2}$ $=\bm{A}_{s}\ @\ \bm{U}_{s}$
$\bm{C}_{2}$ $=\bm{C}_{2}+\bm{L}_{s}^{-}\ @\ \bm{C}_{1}.$
The above sequence of matrix operations has two advantageous properties with respect to data movements. The first two steps of the above sequence share the left matrix operand $\bm{A}$, allowing us to load $\bm{A}$ only once in $\mathtt{L0A}$. Moreover, the third step effectively utilizes the accumulation buffer of the cube unit since $\bm{C}_{2}$ is reused in the last two steps. Once the local scan of a tile of size $\ell$ is computed, a single vector core adds the last value of the previous scanned tile to the current tile (see Lines 141614-16 of Algorithm 2).
1:procedure ScanUL1($\bm{x}$, $s$)
2:  Let $\bm{y}$ be the output array
3:  $\textit{partial}\leftarrow 0$ $\triangleright$ Accumulation value (Vector Unit)
4:  Load $\bm{U}_{s},\bm{L}_{s}^{-},\bm{1}_{s}$ in $\mathtt{L1}$
5:  for each $s^{2}$-tile of $\bm{x}$: $\bm{x}_{\ell}$ do $\triangleright$ Pipelined exec.
6:   Load $\bm{x}_{\ell}$ to $\mathtt{L0A}$ and $\bm{1}_{s}$ to $\mathtt{L0B}$ $\triangleright$ Cube unit
7:   $\bm{C}_{1}\leftarrow\bm{A}_{s}\ @\ \bm{1}_{s}$
8:   Copy $\bm{C}_{1}$ from $\mathtt{L0C}$ to $\mathtt{L1}$
9:   Load $\bm{U}_{s}$ to $\mathtt{L0B}$
10:   $\bm{C}_{2}\leftarrow\bm{A}_{s}\ @\ \bm{U}_{s}$
11:   Load $\bm{L}_{s}^{-}$ in $\mathtt{L0A}$ and $\bm{C}_{1}$ in $\mathtt{L0B}$
12:   $\bm{C}_{2}\leftarrow\bm{C}_{2}+\bm{L}_{s}^{-}\ @\ \bm{C}_{1}$
13:   Copy $\bm{C}_{2}$ from $\mathtt{L0C}$ to $\bm{y}_{\ell}$ in $\mathtt{GM}$
14:   Vector unit waits tile from Cube unit
15:   Copy $\bm{y}_{\ell}$ from $\mathtt{GM}$ to $\mathtt{UB}$ $\triangleright$ Vector unit
16:   $\bm{y}_{\ell}\leftarrow$ $\bm{y}_{\ell}+\textit{partial}$
17:   $\textit{partial}\leftarrow$ last entry of $\bm{y}_{\ell}$
18:   Copy $\bm{y}_{\ell}$ from $\mathtt{UB}$ to $\mathtt{GM}$
19:  Return $\bm{y}$
Algorithm 2 ScanUL1 is adaptation of [13, Alg. 6]
Does cube utilization imply performance?